Solve each system of equation by substitution or addition? x+y=14, 4x-y=-14 show work.
Answer by stanbon(75887)

x+y=14
4x-y=-14
----
Add and solve for "x":
5x = 0
x = 0
------
Substitute and solve for "y":
x +y = 14
0 + y = 14
y = 14
===============
Cheers,
Stan H.